A cyclotron is used to produce a beam of high-energy deuterons that then collide with a target to produce radioactive isotopes for a medical procedure. Deuterons are nuclei of deuterium, an isotope of hydrogen, consisting of one neutron and one proton, with total mass 3.34
Answer #1

a) KE = 1/2 mv^2

5.90E6*1.6E-19 = 0.5*3.34E-27*v^2

v=2.38E7 m/s

b)

q v B = mv^2/r

r = m v/ q B

d = 2 r = 2 m v/q B = 2*3.34e-27*2.38E7/(1.6E-19*1.25)= 0.795 m

c)

I = q/t = N e/t
Not sure if that is A or uA, which is it?
if uA

N/t = I/e = 370E-6/1.6E-19= 2.31E15

if A
N/T = 370/1.6e-19 = 2.31E21